# Potential and Limitation of High-Frequency Cores and Caches

### v1

@@ IV-C New Constraints for Leveraging Improved Performance from Increased Clock Frequency
In order to leverage the improved performance resulting from increased clock frequency in a superconducting environment, subsystems must meet new constraints. One such constraint is the cache bandwidth required to keep up with the faster core. We evaluated the cache bandwidths required for the L1I cache, L1D cache, L2 cache, and L3 cache for the small and large workloads.
@@
Figure 5(a) shows the L1D cache bandwidth for the small workloads. The bars represent different configurations for different small workloads. Figure 5(b) shows the L1D cache bandwidth for the large workloads. The bars represent different configurations for different SPEC2006 workloads. The configurations in this figure are the CryoCore and CryoCache configuration, the SuperCore and CryoCache configuration, and their in-order variants. We chose these because they are more “realistic” than having a configuration with a superconducting cache, since memory systems for superconducting are not as scalable yet [24]. We notice that the L1D cache bandwidth is higher for the SuperCore and CryoCache configuration compared to the CryoCore and CryoCache configuration. For the small workloads, the maximum bandwidth required for the out-of-order architecture is 800 GB/s for the bt workload in the SuperCore and CryoCache configuration, and for in-order architecture, it is 200 GB/s for the bt workload in the in-order SuperCore and CryoCache configuration. For the large workloads, the maximum bandwidth required for the out-of-order architecture is 500 GB/s for the 456.hmmer workload in the SuperCore and CryoCache configuration, and for in-order architecture, it is 130 GB/s for the 456.hmmer workload in the in-order SuperCore and CryoCache configuration.
@@
We observe higher L1D cache bandwidth for the SuperCore and CryoCache configuration compared to the CryoCore and CryoCache configuration, for both the small and large workloads, for both the out-of-order and in-order cores. The higher clock frequency of the core in the SuperCore configuration results in more requests to the L1D cache, which increases the number of accesses per unit time, and hence the bandwidth. The L1I caches, L2 caches, and L3 cache bandwidths are also higher for the SuperCore and CryoCache configuration compared to the CryoCore and CryoCache configuration, for both the small and large workloads, for a similar reason. For the small workloads, the maximum bandwidth required for the L1I cache is 190 GB/s for the tc workload in the SuperCore and CryoCache configuration; for the L2 cache, it is 390 GB/s for the ft workload in the SuperCore and CryoCache configuration; and for the L3 cache, it is 190 GB/s for the cg workload in the SuperCore and CryoCache configuration. For the large workloads, the maximum bandwidth required for the L1I cache is 130 GB/s for the 400.perlbench workload in the SuperCore and CryoCache configuration; for the L2 cache, it is 120 GB/s for the 435.gromacs workload in the SuperCore and CryoCache configuration; and for the L3 cache, it is 70 GB/s for the 470.lbm workload in the SuperCore and CryoCache configuration. The workloads with the highest speedups, both in the small and large workloads, have the highest L1D cache bandwidths. If these bandwidths are not met, the speedups would not be actualized. Therefore, these high bandwidth requirements are a potential limitation of the system, and require redesigning the caches to keep up with the faster core.
@@
Takeaways The workloads with the highest speedups have the highest L1D cache bandwidths. The overall required cache bandwidths are in the range of 130 GB/s to 800 GB/s for the small workloads and 70 GB/s to 500 GB/s for the large workloads. If these bandwidths are not met, the speedups would not be actualized. Therefore, these high bandwidth requirements are a potential limitation of the system, and require redesigning the caches to keep up with the faster core.
V Related Work
@@
Cryogenic computing has been a topic of interest for researchers for a long time. Our work is based on the work of Byun et al. [1] and Min et al. [14], who have proposed a cryogenic core and a cryogenic cache, respectively. Lee et al. have also designed a cryogenic memory [25], which could be used in addition to the cryogenic core and cache to create a complete cryogenic computing system. In this work, we have adapted their designs to gem5 and conducted simulations to evaluate the performance of workloads in cryogenic computing environments.
@@
There has also been some work on superconducting components, part of a larger system, that may be in a room-temperature or cryogenic environment. Herr et al. have proposed a design of a 4-bit RSFQ multiplier-accumulator [29]. They ran tests on its speed and power consumption, and found that was clocked at 11 GHz and consumed 0.2 mW. Zha et al. have proposed a superconducting perceptron-based branch predictor [30]. Nagaoka et al. have proposed a design of a bit-parallel multiplier based on RSFQ technology, which can clock at 52 GHz [31]. Nagaoka et al. also proposed a design of a gate-leveled SFQ multiplier, which can clock at 48 GHz [32]. Obata et al. have proposed an SFQ integer multiplier with a systolic array architecture, clocked at 25 GHz [33]. In this work, we just look at the an entire core or cache in a superconducting environment, and its impact on the performance of workloads. We do not look at the performance of individual components, like a multiplier or a branch predictor, in a superconducting environment in gem5.
There has also been some work on temporal logic in superconducting circuits. Tzimpragos et al. [34, 35] propose that superconducting circuits can compute over temporal relationships between pulse arrivals. They propose the computational relationships between those pulse arrivals can be formalized through an extension to a temporal predicate logic. In this work, we do not look at the temporal logic in superconducting circuits, but rather the high-level implications of a superconducting core and cache on the performance of workloads in a superconducting environment in gem5.
VI Conclusion
In this paper, we presented our methodology for conducting simulations of workloads in cryogenic computing environments using gem5. We then presented the results of our experiments and discussed their implications for the development of cryogenic computing technologies.
Our experiments show that gem5 can be used to simulate workloads in cryogenic computing environments. We used this to show that the performance of the workloads does improve with increasing CryoCore and CryoCache clock frequencies, and characterize the workloads that do not show a substantial improvement in performance. We also showed that the caches would need to be able to provide a higher bandwidth to the core to achieve the speedups shown in the experiments, and provided the bandwidths required to achieve these speedups.
@@ Acknowledgments
We would like to thank the members of the Davis Architecture Research Group (DArchR) for their valuable feedback and support throughout this project. Particularly, we would like to thank Zhantong Qiu for her help in setting up the experiments related to the SimPoints of the large workloads, Harshil Patel for his help in creating the suite of small workloads used in the experiments, and Mahyar Samani for providing feedback on initial drafts of this paper. This work was funded by the Department of Energy under grant number DE-SC0024502.

### v2

@@ IV-C New Constraints for Leveraging Improved Performance from Increased Clock Frequency
In order to leverage the improved performance resulting from increased clock frequency in a superconductor environment, subsystems must meet new constraints. One such constraint is the cache bandwidth required to keep up with the faster core. We evaluated the cache bandwidths required for the L1I cache, L1D cache, L2 cache, and L3 cache for the small and large workloads.
@@
Figure 5(a) shows the L1D cache bandwidth for the small workloads. The bars represent different configurations for different small workloads. Figure 5(b) shows the L1D cache bandwidth for the large workloads. The bars represent different configurations for different SPEC2006 workloads. The configurations in this figure are the CryoCore and CryoCache configuration, the SuperCore and CryoCache configuration, and their in-order variants. We chose these because they are more “realistic” than having a configuration with a superconductor cache, since memory systems for superconductor are not as scalable yet [25]. We notice that the L1D cache bandwidth is higher for the SuperCore and CryoCache configuration compared to the CryoCore and CryoCache configuration. For the small workloads, the maximum bandwidth required for the out-of-order architecture is 800 GB/s for the bt workload in the SuperCore and CryoCache configuration, and for in-order architecture, it is 200 GB/s for the bt workload in the in-order SuperCore and CryoCache configuration. For the large workloads, the maximum bandwidth required for the out-of-order architecture is 500 GB/s for the 456.hmmer workload in the SuperCore and CryoCache configuration, and for in-order architecture, it is 130 GB/s for the 456.hmmer workload in the in-order SuperCore and CryoCache configuration.
@@
We observe higher L1D cache bandwidth for the SuperCore and CryoCache configuration compared to the CryoCore and CryoCache configuration, for both the small and large workloads, for both the out-of-order and in-order cores. The higher clock frequency of the core in the SuperCore configuration results in more requests to the L1D cache, which increases the number of accesses per unit time, and hence the bandwidth. The L1I caches, L2 caches, and L3 cache bandwidths are also higher for the SuperCore and CryoCache configuration compared to the CryoCore and CryoCache configuration, for both the small and large workloads, for a similar reason. For the small workloads, the maximum bandwidth required for the L1I cache is 190 GB/s for the tc workload in the SuperCore and CryoCache configuration; for the L2 cache, it is 390 GB/s for the ft workload in the SuperCore and CryoCache configuration; and for the L3 cache, it is 190 GB/s for the cg workload in the SuperCore and CryoCache configuration. For the large workloads, the maximum bandwidth required for the L1I cache is 130 GB/s for the 400.perlbench workload in the SuperCore and CryoCache configuration; for the L2 cache, it is 120 GB/s for the 435.gromacs workload in the SuperCore and CryoCache configuration; and for the L3 cache, it is 70 GB/s for the 470.lbm workload in the SuperCore and CryoCache configuration. The workloads with the highest speedups, both in the small and large workloads, have the highest L1D cache bandwidths. If these bandwidths are not met, the performance improvement would not be actualized. Therefore, these high bandwidth requirements are a potential limitation of the system, and require redesigning the caches to keep up with the faster core.
@@
Takeaways The workloads with the highest speedups have the highest L1D cache bandwidths. The overall required cache bandwidths are in the range of 130 GB/s to 800 GB/s for the small workloads and 70 GB/s to 500 GB/s for the large workloads. If these bandwidths are not met, the performance improvement would not be actualized. Therefore, these high bandwidth requirements are a potential limitation of the system, and require redesigning the caches to keep up with the faster core.
V Limitations
We only observed the effects of high-level implications of superconductor components and cryogenic semiconductor components on the performance of workloads, i.e., the higher clock frequency of the components in these environments. We did not consider superconductor components having a greater pipeline depth [26] nor temporal logic in superconductor circuits [27, 28], which has an impact on the performance of the workloads. We also did not model the SERDES (serializer-deserializer) circuits that would be required to interface the superconductor components with the room-temperature components, which would have an impact on the performance of the workloads. Instead, we assumed that the interconnect is unchanged from CMOS.
The SuperCore and SuperCache configuration is not feasible in practice. Our paper is a thought experiment to see the potential speedups that could be achieved if an ideal 100 GHz superconductor core and cache were used. In practice, the memory systems for superconductor are not as scalable yet [25], and the bandwidth requirements limit the performance of the workloads.
VI Related Work
@@
Cryogenic semiconductor computing has been a topic of interest for researchers for a long time. Our work is based on the work of Byun et al. [1] and Min et al. [15], who have proposed a cryogenic core and a cryogenic cache, respectively. Lee et al. have also designed a cryogenic memory [29], which could be used in addition to the cryogenic core and cache to create a complete cryogenic semiconductor computing system. In this work, we have adapted their designs to gem5 and conducted simulations to evaluate the performance of workloads in cryogenic semiconductor computing environments.
@@
There has also been some work on superconductor components, part of a larger system, that may be in a room-temperature or cryogenic environment. Herr et al. have proposed a design of a 4-bit RSFQ multiplier-accumulator [33]. They ran tests on its speed and power consumption, and found that was clocked at 11 GHz and consumed 0.2 mW. Zha et al. have proposed a superconductor perceptron-based branch predictor [26]. Nagaoka et al. have proposed a design of a bit-parallel multiplier based on RSFQ technology, which can clock at 52 GHz [34]. Nagaoka et al. also proposed a design of a gate-leveled SFQ multiplier, which can clock at 48 GHz [35]. Obata et al. have proposed an SFQ integer multiplier with a systolic array architecture, clocked at 25 GHz [36]. In this work, we just look at the an entire core or cache in a superconductor environment, and its impact on the performance of workloads. We do not look at the performance of individual components, like a multiplier or a branch predictor, in a superconductor environment in gem5.
There has also been some work on temporal logic in superconductor circuits. Tzimpragos et al. [27, 28] propose that superconductor circuits can compute over temporal relationships between pulse arrivals. They propose the computational relationships between those pulse arrivals can be formalized through an extension to a temporal predicate logic. In this work, we do not look at the temporal logic in superconductor circuits, but rather the high-level implications of a superconductor core and cache on the performance of workloads in a superconductor environment in gem5.
VII Conclusion
In this paper, we presented our methodology for conducting simulations of workloads in cryogenic semiconductor computing environments using gem5. We then presented the results of our experiments and discussed their implications for the development of cryogenic semiconductor computing technologies.
Our experiments show that gem5 can be used to simulate workloads in cryogenic semiconductor computing environments. We used this to show that the performance of the workloads does improve with increasing CryoCore and CryoCache clock frequencies, and characterize the workloads that do not show a substantial improvement in performance. We also showed that the caches would need to be able to provide a higher bandwidth to the core to achieve the speedups shown in the experiments, and provided the bandwidths required to achieve these speedups.
@@ Acknowledgments
We would like to thank the members of the Davis Architecture Research Group (DArchR) for their valuable feedback and support throughout this project. Particularly, we would like to thank Zhantong Qiu for her help in setting up the experiments related to the SimPoints of the large workloads, Harshil Patel for his help in creating the suite of small workloads used in the experiments, and Mahyar Samani for providing feedback on initial drafts of this paper. We would also like to thank D. Scott Holmes and the members of the IEEE IRDS CEQIP team for their valuable feedback on an initial version of this paper. This work was funded by the Department of Energy under grant number DE-SC0024502.